OpenStudy (anonymous):

if cos x = 4over11, whats cos 2A

OpenStudy (anonymous):

\[\cos(x)=\frac{4}{11}\] \[\cos(2x) = ?\]

OpenStudy (anonymous):

8/22 ?

OpenStudy (anonymous):

no, that's just multiplying the numerator and denominator by 2.

OpenStudy (anonymous):

You would have to solve the angle which, when put into the cos function, gives you a value of 4/11

OpenStudy (anonymous):

i dont know how to do that

zepdrix (zepdrix):

One approach would be to use the Cosine Double angle formula. \[\cos 2x = \cos^2 x - \sin^2 x\] We already know the value of cos x, so we can throw that into the square thing... For sine we'll have to set up a triangle.

zepdrix (zepdrix):

Or we could use the other cosine double angle formula actually :D that's probably easier. \[\cos 2x=2\cos^2 x -1\]

zepdrix (zepdrix):

\[2\left(\frac{ 4 }{ 11 }\right)^2-1\]

OpenStudy (anonymous):

You can do this through inverse trig functions which can be thought of as the 'reverse' of trig functions: \[\cos(x) =(\frac{4}{11})\] \[\cancel{\cos^{-1}\cos}(x) = \cos^{-1}(\frac{4}{11})\]

OpenStudy (anonymous):

\[x = \cos^{-1}(\frac{4}{11})\]

zepdrix (zepdrix):

Yah if you are allowed to use a calculator, then math's method is much better :D

OpenStudy (anonymous):

@zepdrix thanks tips

OpenStudy (anonymous):

wait so i just slove the equation

OpenStudy (anonymous):

im still lost

OpenStudy (anonymous):

The question can read: Some angle, when placed into the cosine function will give you a value of 4/11. What value do you get if you put twice that angle into the cos function. 1. first you must solve the angle, through inverse trig functions. 2. then you would multiply that angle by 2 and place it back into the cos function

OpenStudy (anonymous):

i dont know how to do that though

OpenStudy (anonymous):

how do i solve for an angle

OpenStudy (anonymous):



OpenStudy (anonymous):